OpenStudy (anonymous):

A recipe for 120 servings requires 3.5 liters of water. Approximately how many liters of water are needed for 400 servings?

OpenStudy (anonymous):

11.7 I think.

OpenStudy (campbell_st):

well use a ratio \[\frac{400}{120} = \frac{x}{3.5}\] solve for x \[x = \frac{400 \times 3.5}{120}\] just evaluate.